How likely is it that a true relationship between Corrected and sales is “managerially irrelevant” – that is, that the actual coefficient is close enough to 0 that it is not relevant to worry about from a managerial point of view? Assume that “around 0” means in the range of [-0.1, 0.1].

Week Sales Direct Mail Corrected
1 $121,230 9800 42047
2 $212,090 21200 9800
3 $99,980 561000 21200
4 $429,780 41300 56100
5 $496,370 26700 41300
6 $316,450 32100 26700
7 $286,980 52900 32100
8 $496,110 73100 52900
9 $389,080 69500 73100
10 $787,350 54400 69500
11 $446,310 8700 54400
12 $389,410 29900 8700
13 $420,040 24300 29900
14 $629,380 42300 24300
15 $419,370 82000 42300
16 $740,070 60200 82000
17 $498,730 48900 60200
18 $621,780 27900 48900
19 $317,620 37600 27900
20 $427,270 68621 37600
Answer #1

- cet sales = x & corected =y. We are finding linear relationship betxfy ean of line of reg. d y on y es y- y = byx (x-x) or= 2048184256 COUCX,Y)= sxy x = COVcx,y) = ба бу 0.59275 by x = Covcxiy 0-06658 az y - byx () = 13599.33 ? 292 becomes Y: 1359